OpenStudy (anonymous):

Can you use the Law of Sines if you are given a = 31, b = 52 and the measure of angle B equals 56 degrees?

OpenStudy (anonymous):

Care to draw a picture to ensure I know how you are defining your sides and angles?

OpenStudy (anonymous):

I'm pretty sure you can because a side, an opposite side and an angle are provided.

hartnn (hartnn):

see the law of sines \(\Large \dfrac{a}{\sin A}=\dfrac{b}{\sin B}\) you can use this law, if you know 3 terms out of the 4 so do you know 3 terms ?

OpenStudy (anonymous):

I believe so..?

hartnn (hartnn):

which 3 terms? and find the 4th term just keep plugging in values in that

OpenStudy (anonymous):

a/sinA=b/sinB=c/sinC?

OpenStudy (anonymous):

4th term?

hartnn (hartnn):

in this! \(\Large \dfrac{a}{\sin A}=\dfrac{b}{\sin B}\) a,b,sin A, sin B are 4 terms...

OpenStudy (anonymous):

OH okay! So I'm not provided 4 terms?

hartnn (hartnn):

but you know 3 of them, so that u can find the 4th!

OpenStudy (anonymous):

THAT'S what I meant! Thank you again! c:

hartnn (hartnn):

so what did u get A as ?

OpenStudy (anonymous):

.49?

hartnn (hartnn):

you mean sin A = 0.49?? thats correct :)

hartnn (hartnn):

find A from there A = arcsin 0.49

OpenStudy (anonymous):

Right I keep for getting arcsin.

OpenStudy (anonymous):

29.34!

hartnn (hartnn):

correct!: )

OpenStudy (anonymous):

Thank you! I have one more problem then I'll be out of your hair! (For now..)